OpenStudy (anonymous):

Find the slope of the line 3x+5y=-15 a.-3/5 b.5/3 c.3/5 d.-5/3

OpenStudy (anonymous):

slope intercept form line equation is y = mx + c. m = slope and c = intercept 3x+5y=-15 5y=-3x-15 divide each side by 5 y =-3/5x -15/5 y = -3/5x - 5 slope m = -3/5 Option a is right choice. get free math homework help at http://www.mathskey.com